OpenStudy (anonymous):

Find the reciprocal of 3-2i?

OpenStudy (nick88888888):

1/(3+2i) multiply top and bottom by 3-2i 1/(3+2i) (3-2i)/(3-2i) = (3-2i)/(9+4) = (3-2i)/13 = (3/13) + (-2/13) i .

OpenStudy (anonymous):

Ok. Got it. Thanks.